# Kernel vs. Kernel: Exploring How the Data Structure Affects Neural Collapse

### v2

@@ 7 Conclusion
In this paper, we presented a kernel-based approach to understanding the role of data in the emergence of the Neural Collapse (NC) phenomenon. By considering a general kernel function, we first formulated the trace expressions for the variability collapse (NC1) of the features of the data samples. By leveraging these results, we provided theoretical and empirical results to showcase that the NTK does not represent more collapsed features than the NNGP for various Gaussian datasets. Next, to capture the feature-learning aspects of finite-width neural networks, we switched to an ‘adaptive kernel’ approach whose state equations (EoS) facilitate the transition of the post-activation kernel beyond the GP limit. Through this “kernel vs. kernel” approach for limiting NNGP, NTK, and adaptive kernels, we showcased a promising direction to analyze the properties of data for which the NC1 behavior of an actual FCN can be understood. Thus addressing the limitations of the unconstrained features based analysis of NC and explaining the empirical observations of [3] on datasets of varying complexity. We believe that future work on analyzing the EoS for multi-layer FCN and convolutional networks [40] can provide further insights into the depthwise reduction of NC1 for deeper networks and provide a framework for analyzing datasets with more complex distributions.
Acknowledgments and Disclosure of Funding
The authors would like to thank Zhengdao Chen for helpful discussions during the preparation of this manuscript. The work of Tom Tirer is supported by the ISF grant No. 1940/23.
@@ Appendix C General Results for NC1 with Kernels
$\displaystyle\begin{split}\mathbb{E}\left[\mathrm{tr}(\boldsymbol{\Sigma}_{B}(\mathbf{H}))\right]&=\mathbb{E}\left[\frac{1}{C}\sum_{c=1}^{C}\frac{1}{n_{c}^{2}}\sum_{i=1}^{n_{c}}\sum_{j=1}^{n_{c}}\mathbf{Q}(x^{c,i},x^{c,j})\right]-\mathbb{E}\left[\frac{1}{N^{2}}\sum_{c=1}^{C}\sum_{c^{\prime}=1}^{C}\sum_{i=1}^{n_{c}}\sum_{j=1}^{n_{c^{\prime}}}\mathbf{Q}(x^{c,i},x^{c^{\prime},j})\right]\\ &=\left[\frac{1}{2}\sum_{c=1}^{2}\frac{1}{n_{c}^{2}}\left(n_{c}(n_{c}-1)V^{(2)}(c)+n_{c}V^{(1)}(c)\right)\right]\\ &\hskip 20.0pt-\frac{1}{N^{2}}\left[\sum_{c=1}^{2}\left(n_{c}(n_{c}-1)V^{(2)}(c)+n_{c}V^{(1)}(c)\right)\right]\\ &\hskip 20.0pt-\frac{1}{N^{2}}\left[2\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}V^{(3)}(c=1,c^{\prime}=2)\right]\\ &=\left[\sum_{c=1}^{2}\left(\frac{1}{2n_{c}^{2}}-\frac{1}{N^{2}}\right)\left(n_{c}(n_{c}-1)V^{(2)}(c)+n_{c}V^{(1)}(c)\right)\right]-\frac{2n_{1}n_{2}}{N^{2}}V^{(3)}(1,2)\end{split}$ (31)
@@
Here, $\Delta_{h.o.t}$ corresponds to higher order terms given by [53]:
@@
where, based on the well-studied concentration of sample covariance matrices around the statistical covariance [54], $\Delta_{h.o.t}$ tend to 00 0 for large $n_{c}$ values.
@@
Based on the standard result on the expectation of ratios [53], we get:
@@ D.2 NC1 of limiting NTK with ReLU activation
The recursive relationship between the NTK and NNGP [31, 35] can be given as follows (13):
@@ E.1.1 Calculating $\mathbb{E}\left[Q_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})\right]$
Thus, based on the expansion of $(1+u)^{-1}=1-u+u^{2}-u^{3}+\cdots$, we obtain the following cleaner approximation of:
@@ E.2.1 Calculating $\mathbb{E}\left[\Theta_{NTK-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})\right]$
Thus, based on the expansion of $(1+u)^{-1}=1-u+u^{2}-u^{3}+\cdots$, we obtain the following cleaner approximation of:
@@ Appendix F Activation Variability Relative to Data
$\bullet$ Addressing misleading $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values: Consider the case where $\sigma_{1},\sigma_{2}\to 0$. Then Theorem 5.1 for $Q_{GP-ReLU}$ indicates that $\mathbb{E}\left[\mathcal{N}\mathcal{C}_{1}(\mathbf{H})\right]\to 0$ (considering smaller fluctuations from $\Delta_{h.o.t}$) in the balanced class setting. Such an observation can be misleading if one were to ignore $\mathcal{N}\mathcal{C}_{1}(\mathbf{X})$. For instance, such an empirical result while training deep neural networks fails to differentiate between settings where the network learned meaningful features and learned to classify complex datasets or was simply able to leverage the already collapsed data vectors. This applies to Erf activation as well. We justify this argument with the following experiment. For a sample size $N$ chosen from $\{128,256,512,1024\}$, and input dimension $d_{0}$ chosen from $\{1,2,8,32,128\}$, we sample the vectors $\mathbf{x}^{1,i}\sim\mathcal{N}(-10*\mathbf{1}_{d_{0}},\mathbf{I}_{d_{0}}),i\in[N/2]$ for class 11 11 and $\mathbf{x}^{2,j}\sim\mathcal{N}(10*\mathbf{1}_{d_{0}},\mathbf{I}_{d_{0}}),j\in[N/2]$ for class 22 22 as our dataset. From Figure 5(a), 5(b), observe that $\mathcal{N}\mathcal{C}_{1}(\mathbf{H}|\mathbf{X})$ values for $Q_{GP-Erf}$ can be orders of magnitude larger than $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$, and for high-dimensions $\mathcal{N}\mathcal{C}_{1}(\mathbf{H}|\mathbf{X})>1$. Essentially, the raw data is ‘more’ collapsed than the activations in these settings. Similar observations can be made for the NTK $\Theta_{NTK-Erf}$ in Figure 5(c), 5(d).
@@
Appendix G Numerical solutions of EoS
We solve the EoS using the Newton-Krylov method with an annealing schedule (as originally proposed by [40]) using the scipy.optimize.newton_krylov python API. We initialize $\mathbf{C}$ with the GP limit value of $(\sigma_{w}^{2}/d_{0})\mathbf{I}_{d_{0}}$ and choose a large annealing factor (ex: $10^{5}$) as the value for $d_{1}$. The result of optimizing with newton_krylov is a new $\mathbf{C}$, which in addition to a lower annealing factor is used as an input for the next newton_krylov function call. This loop is repeated until the end of an annealing schedule. For instance, to analyze the EoS corresponding to $d_{1}=500$, we choose the following list of step-wise annealing factors:
$\bullet$ Comparing the spectrum of weight covariance matrices: Since $\mathbf{C}$ is subject to change while obtaining the stable state of the EoS, we analyze its initial and final (normalized) spectra for two different datasets of dimension $d_{0}=32$ and $N=1024$. Dataset 1: $\mathbf{x}^{1,i}\sim\mathcal{N}(-2*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),i\in[N/2]$, $\mathbf{x}^{2,j}\sim\mathcal{N}(2*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),j\in[N/2]$. Dataset 2: $\mathbf{x}^{c,i}\sim\mathcal{N}(\mathbf{0}_{d_{0}},4*\mathbf{I}_{d_{0}}),i\in[N/2],c\in[2]$ The first dataset is our running example, and the second is pure random noise data. Surprisingly, we observed that the EoS solution captures correlations in the data for both datasets, which is reflected in its final spectrum. In particular, the singular values shift from being constant at initialization to exhibiting a decay in their values (Figure 7). Such a shift does not exactly match the case of 22 22L-FCN because of (1) the difference in the dynamics of GD and Newton-Krylov with annealing, and (2) we start with a GP-based initial value for $\mathbf{C}$ in EoS. A rigorous analysis of the EoS dynamics is an open research direction (as also highlighted by the Seroussi et al. [40]). Nonetheless, the EoS offers a richer data-dependent setup to analyze the activations and weights, than the UFM.
(a) EoS: Dataset 1
(b) 2L-FCN: Dataset 1
(c) EoS: Dataset 2
(d) 2L-FCN: Dataset 2
Figure 7: Normalized singular values sorted in descending order $\lambda_{i}/\lambda_{max},\forall i\in[32]$ for $\mathbf{C}\in\mathbb{R}^{d_{0}\times d_{0}}$ in case of EoS and for $\mathbf{W}^{(1)\top}\mathbf{W}^{(1)}\in\mathbb{R}^{d_{0}\times d_{0}}$ in case of 2L-FCN. Here init represents the initialized state of an EoS and 2L-FCN in their respective plots. The final state of EoS is obtained by solving it using Newton-Krylov. The 2L-FCN with $d_{1}=500$ is trained for $10,000$ epochs using Gradient Descent with a learning rate $10^{-3}$, weight decay $10^{-6}$, $\sigma_{w}=1$ and $\sigma_{b}=0$. The EoS follows an annealing schedule with a final factor 500500 500500.
Appendix H Additional Experiments
Compute Resources: All the experiments in this paper were executed on a machine with 1616 1616 GB of host memory and 88 88 CPU cores. Experiments with the EoS on datasets of varying dimensions and sample sizes took the longest time $\approx 1$ hour to finish.
@@ Appendix H Additional Experiments
(c) $Q_{GP-ReLU}$
(d) $\Theta_{NTK-ReLU}$
Figure 8: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the post-activation NNGP kernel ($Q^{(1)}_{GP}$) and NTK ($\Theta^{(2)}$) corresponding to Erf, ReLU activations. The dimension $d_{0}$ on the x-axis is chosen from $\{1,2,8,32,128\}$. For $(n_{1},n_{2})$ such that $n_{1}+n_{2}=N=1024$, we sample the vectors $\mathbf{x}^{1,i}\sim\mathcal{N}(-2*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{1,i}=-1,i\in[n_{1}]$ for class 11 11 and $\mathbf{x}^{2,j}\sim\mathcal{N}(2*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{2,j}=1,j\in[n_{2}]$ for class 22 22 as our dataset.
@@
Figure 10: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the limiting kernels, adaptive kernel (EoS) with final annealing factor $d_{1}=500$ and 2L-FCN with $d_{1}=500$ and Erf activation. The dimension $d_{0}$ on the x-axis is chosen from $\{1,2,8,32,128\}$. For a tuple $n_{c}=(n_{1},n_{2},n_{3},n_{4})$ such that $n_{1}+n_{2}+n_{3}+n_{4}=N=1024$, we sample the vectors $\mathbf{x}^{1,i}\sim\mathcal{N}(-6*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{1,i}=-3,i\in[n_{1}]$ for class 11 11, $\mathbf{x}^{2,j}\sim\mathcal{N}(-2*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{2,j}=-1j\in[n_{2}]$ for class 22 22, $\mathbf{x}^{3,k}\sim\mathcal{N}(2*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{3,k}=1,k\in[n_{3}]$ for class 33 33 and $\mathbf{x}^{4,l}\sim\mathcal{N}(6*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{4,l}=3,l\in[n_{4}]$ for class 44 44.
(a) $Q_{GP-Erf}$
(b) $\Theta_{NTK-Erf}$
(c) EoS
(d) 2L-FCN
(a) $Q_{GP-Erf}$
(b) $\Theta_{NTK-Erf}$
(c) EoS
(d) 2L-FCN
Figure 12: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the limiting kernels, adaptive kernel (EoS) with final annealing factor $d_{1}=500$ and 2L-FCN with $d_{1}=500$ and Erf activation. The dimension $d_{0}$ on the x-axis is chosen from $\{1,2,8,32,128\}$. For a particular $N$, we sample the vectors $\mathbf{x}^{1,i}\sim\mathcal{N}(-2*\mathbf{1}_{d_{0}},\mathbf{I}_{d_{0}}),y^{1,i}=-1,i\in[N/2]$ for class 11 11 and $\mathbf{x}^{2,j}\sim\mathcal{N}(2*\mathbf{1}_{d_{0}},\mathbf{I}_{d_{0}}),y^{2,j}=1,j\in[N/2]$ for class 22 22.
(a) $Q_{GP-Erf}$
(b) $\Theta_{NTK-Erf}$
(c) EoS
(d) 2L-FCN
(a) $L=3$
(b) $L=4$
(c) $L=5$
(d) $L=6$
Appendix I Impact Statement
This paper aims to address the limitations of the unconstrained features model to understand the role of data on the Neural Collapse phenomenon. Our work does not have any direct negative societal impact. On the contrary, our work takes a step forward in understanding the characteristics of datasets that govern the performance of deep neural network classifiers. Thus, laying the groundwork for theoretically informed training practices for wide societal use.
Appendix J Limitations and Future Work
In certain cases, we have observed that none of the kernel methods approximate the 2L-FCN reasonably. One such instance is the following, where we sample $\mathbf{x}^{1,i}\sim\mathcal{N}(-2*\mathbf{1}_{d_{0}},4*\mathbf{I}_{d_{0}}),y^{1,i}=-1,i\in[N/2]$ for class 11 11 and $\mathbf{x}^{2,j}\sim\mathcal{N}(2*\mathbf{1}_{d_{0}},4*\mathbf{I}_{d_{0}}),y^{2,j}=1,j\in[N/2]$ for class 22 22 of our dataset. Essentially, these are scenarios where there is a significant overlap between samples of the two classes. First, we note that we had to increase the learning rate of our 2L-FCN from $10^{-3}$ to $5\cdot 10^{-3}$ and run GD for 20002000 20002000 epochs for convergence. For dimensions $d_{0}=\{8,16,32\}$, the EoS reasonably approximates the 2L-FCN but for $d_{0}=\{64,128\}$, the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values for 2L-FCN turned out to be almost twice as large as the EoS (see Figure 13). To this end, we leave modifications to the EoS for handling such noisy data cases and different activation functions as future work.
Additionally, we highlight the difficulties in the theoretical/empirical analysis of NC1 with EoS. The primary bottleneck is a lack of rigorous study on the existence and uniqueness of solutions (As also highlighted by [40]). Since we deviate from the lazy regime and deal with kernels in the feature learning setup, we cannot expect simpler closed-form solutions like the limiting NNGP/NTK for the EoS. However, analytical solutions to the EoS can sometimes be time-consuming and require a manual selection of the annealing schedule. This is a tradeoff that can be improved with future research. Furthermore, the role of scaling $N,d_{0},d_{1}$ on NC1 is yet to be fully understood and we hope that our analysis lays the groundwork for such efforts.

### v3

@@ 8 Conclusion
In this paper, we explored whether kernel-based approaches can help understand the role of data in the emergence of the NC phenomenon. By considering a general kernel function, we first formulated the trace expressions for the variability collapse (NC1) of the features of the data samples. By leveraging these results, we provided theoretical and empirical results to showcase that the NTK does not represent more collapsed features than the NNGP for various Gaussian datasets. During this investigation we showed how the choice of the nonlinear activation function affects NC1. Next, to capture the feature-learning aspects of finite-width neural networks, we switched to a data-aware GP kernel approach whose state equations (EoS) facilitate the transition of the post-activation kernel beyond the NNGP limit. We empirically showed that it yields lower NC1 than NNGP but may not be aligned with trends of FCNs. The key message of the paper is that advancements in modeling NNs by adaptive kernels can benefit NC analysis and lead to a considerable leap compared to current UFM based theory.
Acknowledgment
TT was supported by ISF grant No. 1940/23 and MOST grant No. 0007091. The authors also thank the anonymous reviewers for their valuable feedback.
@@ Appendix C General Results for NC1 with Kernels
$\displaystyle\begin{split}\mathbb{E}\left[\mathrm{tr}(\boldsymbol{\Sigma}_{B}(\mathbf{H}))\right]&=\mathbb{E}\left[\frac{1}{C}\sum_{c=1}^{C}\frac{1}{n_{c}^{2}}\sum_{i=1}^{n_{c}}\sum_{j=1}^{n_{c}}\mathbf{Q}(x^{c,i},x^{c,j})\right]-\mathbb{E}\left[\frac{1}{N^{2}}\sum_{c=1}^{C}\sum_{c^{\prime}=1}^{C}\sum_{i=1}^{n_{c}}\sum_{j=1}^{n_{c^{\prime}}}\mathbf{Q}(x^{c,i},x^{c^{\prime},j})\right]\\ &=\left[\frac{1}{2}\sum_{c=1}^{2}\frac{1}{n_{c}^{2}}\left(n_{c}(n_{c}-1)V^{(2)}(c)+n_{c}V^{(1)}(c)\right)\right]\\ &\hskip 20.0pt-\frac{1}{N^{2}}\left[\sum_{c=1}^{2}\left(n_{c}(n_{c}-1)V^{(2)}(c)+n_{c}V^{(1)}(c)\right)\right]-\frac{1}{N^{2}}\left[2\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}V^{(3)}(c=1,c^{\prime}=2)\right]\\ &=\left[\sum_{c=1}^{2}\left(\frac{1}{2n_{c}^{2}}-\frac{1}{N^{2}}\right)\left(n_{c}(n_{c}-1)V^{(2)}(c)+n_{c}V^{(1)}(c)\right)\right]-\frac{2n_{1}n_{2}}{N^{2}}V^{(3)}(1,2)\end{split}$ (31)
@@
Here, $\Delta_{h.o.t}$ corresponds to higher order terms given by (Seltman, 2012):
@@
where, based on the well-studied concentration of sample covariance matrices around the statistical covariance (Vershynin, 2012), $\Delta_{h.o.t}$ tend to 00 0 for large $n_{c}$ values.
@@
Based on the standard result on the expectation of ratios (Seltman, 2012), we get:
@@ D.2 NC1 of limiting NTK with ReLU activation
The recursive relationship between the NTK and NNGP (Lee et al., 2019; Tirer et al., 2022) can be given as follows (13):
@@ E.1.1 Calculating $\mathbb{E}\left[Q_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})\right]$
Thus, based on the expansion of $(1+u)^{-1}=1-u+u^{2}-u^{3}+\cdots$, we obtain the following cleaner approximation:
@@ E.2.1 Calculating $\mathbb{E}\left[\Theta_{NTK-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})\right]$
Thus, based on the expansion of $(1+u)^{-1}=1-u+u^{2}-u^{3}+\cdots$, we obtain the following cleaner approximation:
@@ Appendix F Activation Variability Relative to Data
$\bullet$ Addressing misleading $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values: Consider the case where $\sigma_{1},\sigma_{2}\to 0$. Then Theorem 5.1 for $Q_{GP-ReLU}$ indicates that $\mathbb{E}\left[\mathcal{N}\mathcal{C}_{1}(\mathbf{H})\right]\to 0$ (considering smaller fluctuations from $\Delta_{h.o.t}$) in the balanced class setting. Such an observation can be misleading if one were to ignore $\mathcal{N}\mathcal{C}_{1}(\mathbf{X})$. For instance, such an empirical result while training deep neural networks fails to differentiate between settings where the network learned meaningful features and learned to classify complex datasets or was simply able to leverage the already collapsed data vectors. This applies to Erf activation as well. We justify this argument with the following experiment. For a sample size $N$ chosen from $\{128,256,512,1024\}$, and input dimension $d_{0}$ chosen from $\{1,2,8,32,128\}$, we sample the vectors $\mathbf{x}^{1,i}\sim\mathcal{N}(-10*\mathbf{1}_{d_{0}},\mathbf{I}_{d_{0}}),i\in[N/2]$ for class 11 11 and $\mathbf{x}^{2,j}\sim\mathcal{N}(10*\mathbf{1}_{d_{0}},\mathbf{I}_{d_{0}}),j\in[N/2]$ for class 22 22 as our dataset. From Figure 6(a), 6(b), observe that $\mathcal{N}\mathcal{C}_{1}(\mathbf{H}|\mathbf{X})$ values for $Q_{GP-Erf}$ can be orders of magnitude larger than $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$, and for high-dimensions $\mathcal{N}\mathcal{C}_{1}(\mathbf{H}|\mathbf{X})>1$. Essentially, the raw data is ‘more’ collapsed than the activations in these settings. Similar observations can be made for the NTK $\Theta_{NTK-Erf}$ in Figure 6(c), 6(d).
@@
Appendix G Training Details and Additional Experiments
Setup: We train a 2L-FCN with $d_{1}=500,\sigma_{w}=1,\sigma_{b}=0$ and Erf activation using (vanilla) Gradient Descent with a learning rate of $10^{-3}$ and weight-decay $10^{-6}$ for 10001000 10001000 steps to reach the terminal phase of training (see Figure 8) for plots with $N=1024$ and vectors $\mathbf{x}^{1,i}\sim\mathcal{N}(-2*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{1,i}=-1,i\in[N/2]$ for class 11 11 and $\mathbf{x}^{2,j}\sim\mathcal{N}(2*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{2,j}=1,j\in[N/2]$ for class 22 22). The ReLU activation experiments use a learning rate of $10^{-4}$. To numerically solve the EoS, we employ the approach described in Section 6.2. All the experiments in this paper were executed on a machine with 1616 1616 GB of host memory and 88 88 CPU cores. Experiments with the EoS on datasets of varying dimensions and sample sizes took the longest time $\approx 1$ hour.
(a) Loss ($d_{0}=1$)
(b) Sign prediction ($d_{0}=1$)
(c) Loss ($d_{0}=8$)
(d) Sign prediction ($d_{0}=8$)
Figure 8: Training loss and predictions (sign) of the 2L-FCN model with $d_{1}=500$ and Erf activation.
(a) $L=3$
(b) $L=4$
(c) $L=5$
(d) $L=6$
On increasing the number of classes with separable data.
Similar to the formulation of $\mathcal{D}_{1}(N,d_{0})$ for $C=2$, we formulate $\mathcal{D}_{2}(N,d_{0})$ for $C=4$ as follows:
The trends in $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ for EoS and 2L-FCN hold even for datasets $\mathcal{D}_{2}(N,d_{0})$ with $C>2$. Observe from Figure 10 that the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values for the kernels $Q_{GP-Erf},\Theta_{NTK-Erf}$, EoS and for 2L-FCN are consistently lower than the $C=2$ case as shown in Figures 2(a), 2(b), 2(c), 4(b), even when comparing them at $d_{0}=1$. We believe this is because $\mathcal{D}_{2}(N,d_{0})$ is constructed by adding 22 22 new classes to $\mathcal{D}_{1}(N,d_{0})$ whose distance between the means is much larger than the within class co-variances (see Figure 11). Nonetheless, similar to the $C=2$ case, the EoS deviates from the NNGP behavior and tends to approximate the trends of 2L-FCN for high dimensional data with a sufficiently large sample size.
@@ Appendix G Training Details and Additional Experiments
(c) EoS
(d) 2L-FCN
Figure 10: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the limiting kernels, adaptive kernel (EoS) with final annealing factor 500500 500500 and 2L-FCN with $d_{1}=500$ and Erf activation on dataset $\mathcal{D}_{2}(N,d_{0})$ (as per (156)).
(a) $Q_{GP-Erf}$
(b) $\Theta_{NTK-Erf}$
(c) EoS
(d) 2L-FCN
Class imbalance hurts EoS in approximating 2L-FCN.
One of the key conditions for the EoS to approximate the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ behavior of a 2L-FCN is the presence of sufficient data samples (Seroussi et al., 2023) (depending on the scale of the input dimension $d_{0}$ as shown above). Thus, extreme class imbalances can lead to biased $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ trends in EoS. To this end, we consider a collection of imbalanced datasets based on $\mathcal{D}_{1}(N,d_{0})$ where $N=2048$ is split into two classes as follows: Case 1: $(768,1280)$, Case 2: $(512,1536)$, Case 3: $(256,1792)$. We observe that as the imbalance ratio increases, the magnitude of $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ in 2L-FCN and kernel methods show opposing trends (see Figure 12). We make a similar observation by considering $\mathcal{D}_{2}(N,d_{0})$ with $C=4$ and $N=1024$ with 3 different imbalance ratios as in Figure 13.
@@
Figure 13: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the limiting kernels, adaptive kernel (EoS) with final annealing factor $d_{1}=500$ and 2L-FCN with $d_{1}=500$ and Erf activation. The dimension $d_{0}$ on the x-axis is chosen from $\{1,2,8,32,128\}$. For a tuple $n_{c}=(n_{1},n_{2},n_{3},n_{4})$ such that $n_{1}+n_{2}+n_{3}+n_{4}=N=1024$, we sample the vectors $\mathbf{x}^{1,i}\sim\mathcal{N}(-6*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{1,i}=-3,i\in[n_{1}]$, $\mathbf{x}^{2,j}\sim\mathcal{N}(-2*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{2,j}=-1j\in[n_{2}]$, $\mathbf{x}^{3,k}\sim\mathcal{N}(2*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{3,k}=1,k\in[n_{3}]$ and $\mathbf{x}^{4,l}\sim\mathcal{N}(6*\mathbf{1}_{d_{0}},0.25*\mathbf{I}_{d_{0}}),y^{4,l}=3,l\in[n_{4}]$.